### v3

# Machine Learning $H$-theorem

Abstract
$H$-theorem provides a microscopic foundation of the Second Law of Thermodynamics and is therefore essential to establishing statistical physics, but at the same time, $H$-theorem has been subject to controversy that in part persists till this day. To better understand $H$-theorem and its relation to the arrow of time, we study the equilibration of randomly oriented and positioned hard disks with periodic boundary conditions. Using a model based on the DeepSets architecture, which imposes permutation invariance of the particle labels, we train a model to capture the irreversibility of the $H$-functional.
I Introduction
The discovery of $H$-theorem by Ludwig Boltzmann in 1872 boltzmann1872waermegleichgewicht played a pivotal role in establishing statistical mechanics as a framework that provides thermodynamics with a microscopic origin and liberates it from its axiomatic foundations. At the same time, $H$-theorem was for a long time the topic of much debate https://doi.org/10.1002/andp.18972960216 ; zermelo2 ; zermelo1 , parts of which persist to this day Villani2003 ; Borsoni2024 ; https://doi.org/10.1002/cpa.10012 ; DiPerna1989 . One lingering issue is the Loschmidt paradox Cercignani2006-if ; Cercignani1988 , which relates to the remarkable feature of $H$-theorem that there exists an $H$-functional which can only decrease in time and therefore signal the presence of an arrow of time eddington1929nature despite the system being microscopically reversible doi:10.1073/pnas.11.7.436 , which means that the particles of which the gas is composed undergo interactions that are time-reversal invariant. This paradox is now better understood, in part because of the discovery of BBGKY hierarchy grad1949 ; ecc82589-c02d-30e9-9891-fe9629a2fd13 , which must be truncated by assuming molecular chaos so that one is left with the Boltzmann equation. The assumption of molecular chaos introduces an arrow of time; while Lanford’s theorem lanford1975 ; AST_1976__40__117_0 ; RevModPhys.52.569 proves its validity in the Boltzmann–Grad limit grad1949 for short times, no general proof exists for arbitrary times or interaction types, so its adoption remains, to some extent, a leap of faith. Nevertheless, the Boltzmann equation has been found to be a valuable tool for deriving continuum equations from first principles that has been very widely used for more than a century Cercignani1988 ; chapman1990mathematical in fields ranging from solid state physics PhysRevB.97.045105 ; mueller2009graphenenearlyperfect ; PhysRevLett.115.216806 to high energy physics De_Groot1980-px ; Cercignani2002-lo ; Mueller_2004 as well as in active matter PhysRevE.83.030901 ; PhysRevE.74.022101 ; grosvenor2025hydrodynamicsboostinvariancekinetictheory ; rubenliermicroscopic .
In recent years, several works tried to exploit the emergence of an arrow of time for many-particle systems as a way to better understand sequential data. For example, in Ref. Seif2021 , the authors show that by putting trajectories and their time-reverses into a binary classifier, it is possible to extract Crooks fluctuation theorem PhysRevE.60.2721 ; PhysRevLett.78.2690 from the predicted probability of the binary classifier. In Ref. kuroyanagi2025deeplearningthermodynamiclaws , the authors use siamese neural network to compare pairs of image sets at different times for a molecular dynamics simulation which involves a moving piston, enabling the authors to construct an entropy. This equilibrium entropy is shown to be consistent with thermodynamic axioms formulated by Lieb and Yngvason LIEB19991 and identical to the entropy of the van der Waals gas up to an affine transformation. Similarly, in Ref. Wei_2018_CVPR , videos played forward and reversed are classified, and the classifier is then used to identify the spatial region in the video that has the strongest arrow of time. Lastly in Ref. rahaman2019learningarrowtime , the authors learn a linear function that separates states early and late in time for games that have an irreversible element in it and therefore require safe exploration moldovan2012safeexplorationmarkovdecision ; inproceedings . The authors show that when such a function is learned for a two-component stochastic process the function has quantitative overlap with entropy up to an affine transformation.
In this work, we learn to output a value that resembles $H$-functional up to an affine transformation for sequential data of a numerical simulation of randomly initialized hard disks PhysRevA.1.18 ; PhysRevE.92.022131 ; PhysRevLett.18.988 ; ORBAN1967620 . We infer the $H$-functional using an approach that is similar to that of Ref. rahaman2019learningarrowtime where a linear function is derived whose cost function is such that the learning processes pushes the model to learn to attribute higher output values to input values corresponding to later times.
The structure of this work is as follows. In Sec. II, we discuss Boltzmann equation and the related $H$-theorem. In Sec. III, we discuss the simulation performed to generate the data for which we wish to learn $H$-theorem, which is randomly initialized hard disks. In Sec. IV, we discuss the learning of $H$-theorem using a model that adheres to permutation invariance of the particle labels using the DeepSets architecture and learns a function that approaches $H$-theorem by requiring that it gives a higher value to input at later times.
II Boltzmann equation
We consider the number of particles $dN$ in a region in phase space to be given by
$dN=f(\mathbf{r},\mathbf{v},t)d\mathbf{r}d\mathbf{v}~~,$ (1)
where $\mathbf{r}$ is the particle position, $\mathbf{v}$ is its velocity and $f(\mathbf{r},\mathbf{v},t)$ is the distribution function. In the presence of a force $\mathbf{F}$, the Boltzmann equation is given by
$\frac{\partial f}{\partial t}+\mathbf{v}\cdot\frac{\partial f}{\partial\mathbf{r}}+\frac{\mathbf{F}}{m}\cdot\frac{\partial f}{\partial\mathbf{v}}=\left(\frac{\partial f}{\partial t}\right)_{\text{self}}~~,$ (2)
where $m$ is the particle mass. In this work, we set $\mathbf{F}=0$ since the interactions are exclusively of the hard-sphere kind, and we also do not account for spatial dependence. The term in the right hand side of (2) is the collision integral for which one can find closure with the molecular chaos assumption Cercignani1988 ; grad1958principles . For the case of hard disks, the collision integral is given by chapman1990mathematical ; Cercignani1988
$\begin{split}&\left(\frac{\partial f}{\partial t}\right)_{\text{self}}=\\ &\iint d\mathbf{v}_{2}db\,u\left(f(\mathbf{v}^{\prime}_{1},t)f(\mathbf{v}^{\prime}_{2},t)-f(\mathbf{v},t)f(\mathbf{v}_{2},t)\right),\end{split}$ (3)
where $b$ is the impact parameter and $\mathbf{v}^{\prime}_{1}$ and $\mathbf{v}^{\prime}_{2}$ are velocities of particles going into the collision which are tied to $\mathbf{v}$ and $\mathbf{v}_{2}$ through energy and momentum balance. Galilean invariance makes it that only the relative velocities $\mathbf{u}=\mathbf{v}-\mathbf{v}_{2}$ and $\mathbf{u}^{\prime}=\mathbf{v}^{\prime}_{1}-\mathbf{v}^{\prime}_{2}$ describe a unique collision, and the scattering angle $\chi$ that gives the relative angles between these two vectors is given by (see Fig. 1)
$b=d\cos(\chi/2)~~,$ (4)
where $d$ is the disk diameter.
Figure 1: Depiction of an elastic collision between two hard disks.
To solve (2) analytically for the exact collision operator is incredibly difficult. One way to do so is perturbatively, using the Chapman-Enskog expansion chapman1990mathematical . There exist however numerical approaches to solving (2) Mouhot_2006 and recently it has been explored whether a neural network can be used to deal with the collision integral Xiao_2021 ; Xiao_2023 . Analytical challenges notwithstanding, $H$-theorem provides us with an excellent handle for understanding how (2) irreversibly moves towards an equilibrium state. To wit, $H$-theorem tells us that there exists an $H$-functional given by
$H(t)=\int d\mathbf{v}f(\mathbf{v},t)\log(f(\mathbf{v},t))~~,$ (5)
for which it holds that upon plugging in (2), one obtains
$\frac{\partial}{\partial t}H(t)\leq 0~~.$ (6a)
Furthermore, in two dimensions the equilibrium state is characterized by
$\begin{split}\frac{\partial}{\partial t}H_{0}(t)&=0~~,\\ f_{0}(\mathbf{v},t)&=\frac{nm}{2\pi{k_{\text{b}}}T}\exp(-\frac{m}{2{k_{\text{b}}}T}|\mathbf{v}-\mathbf{v}_{0}|^{2})~~,\end{split}$ (6b)
where $n$ is the particle density, $T$ is temperature and $\mathbf{v}_{0}$ is the average velocity. $f_{0}(\mathbf{v},t)$ is called the Maxwell-Boltzmann distribution. The striking elegance of (6) which emerges from the chaotic nature of hard disk collisions tells us that something very simple can be learned from very chaotic data and thereby underpins our machine learning approach.
III Hard disk simulation
Let us now specify the simulation. The simulation is based on code from hill2020mb2d and is initialized with 1000 hard disks of equal mass and diameter with random orientation and position in a square box with periodic boundary conditions. The initial velocity magnitude is chosen to be uniform, as a uniformly distributed velocity magnitude gives rise to an $H$-functional that undergoes only very minor changes through time and therefore it is not easy to see the downward trend dictated by $H$-theorem. For every timestep, it is checked whether two particles overlap, in which case an elastic collision takes place. This collision is constrained by energy and momentum balance and the requirement that momentum is only transferred in the direction parallel to the line of disk centers (see Fig. 1). Since all particles have the same mass, these collision rules can be described in terms of the simple relations
$\mathbf{n}\cdot\mathbf{u}^{\prime}=-\mathbf{n}\cdot\mathbf{u}~~,~~\mathbf{z}\times\mathbf{n}\times\mathbf{u}^{\prime}=\mathbf{z}\times\mathbf{n}\times\mathbf{u}~~,$ (7)
where $\mathbf{n}$ is the line of disk centers and $\mathbf{z}$ is the out of plane direction. This collision only takes place if the relative velocity points oppositely to the line of disk centers, i.e. when it holds that
$\mathbf{u}\cdot\mathbf{n}<0~~.$ (8)
as otherwise the particles are moving away from each other. If this constraint is not imposed, particles will get stuck together. The diameter of the disks is taken to be $d=6\cdot 10^{-3}$ whereas the periodic box has a unit surface. When the frame rate is too small, it is possible that certain collisions are missed. In Fig. 2, we show that at a frame rate of around $60$, the number of collisions starts to plateau, so this is the frame rate we take.
Figure 2: Number of collisions that are observed in a single run averaged over 5 runs vs the frame rate. The frame rate is the inverse of the time step dtdt of the simulation.
All velocities are saved for each timestep giving rise to an amount of $N_{t}=500$ time steps, and we considered $N_{\text{run}}=500$ unique runs. In Fig. 3, we show the distribution of the hard disks and compare it to the Maxwell-Boltzmann distribution of (6b).
Figure 3: Picture of theoretically predicted Maxwell-Boltzmann distribution together with average and final frame of hard disk simulation for the four different runs that are also considered in Fig. 5. For constructing the histograms, we used bin size 0.005. The uniform velocity magnitude at which the disks are initialized is given by $v=0.12$, which means the Maxwell-Boltzmann distribution whose expression is given by (6b) has temperature ${k_{\text{b}}}T=0.0076$ and mass $m=1$. The averaging started after 20% of the run was completed.
IV Machine learning $H$-theorem
Let us now describe the machine learning process. At every training step, we have as input sequential data $\mathbf{V}^{[n]}_{t}$ for all $t$ and $n$, where $t$ corresponds to the time step and $n$ corresponds to the run. For any run $n$, we have
$\mathbf{v}^{(i),[n]}_{t}\in\mathbf{V}^{[n]}_{t}~~$, (9)
where $\mathbf{v}^{(i),[n]}_{t}$ is the two-dimensional velocity vector of the $i$th particle. In order to effectively learn the $H$-functional, we utilize the permutation invariants of the particle labels, which is done with the DeepSets architecture zaheer2018deepsets . In particular, we take the model to be of the form
[…]
where $\rho$ and $\phi$ encode a neural network with trainable weights that treats any value of run $n$, any run $t$ and any particle $i$ on equal footing. Following rahaman2019learningarrowtime , we consider a loss function per timestep $t\leq N_{t}-1$ and run $n$ that is given by
$\begin{split}&\mathcal{L}^{(0)}\left(\mathbf{V}^{[n]}_{t},\mathbf{V}^{[n]}_{t+1}\right)=h\left(\mathbf{V}^{[n]}_{t}\right)-h\left(\mathbf{V}^{[n]}_{t+1}\right)\\ &+\lambda\,\!\left(h\left(\mathbf{V}^{[n]}_{t+1}\right)-h\left(\mathbf{V}^{[n]}_{t}\right)\right)^{2}\end{split}$ (10b)
The loss of (10b) contains a term that pushes to learn a function consistent with the monotonicity property (6a) as well as the term with coefficient $\lambda$ that serves as a regularization. In Fig. 4, we give a graphical representation of the learning system of (10) and we see that the pairwise comparative nature of (10b) makes it so that this learning system effectively turns into a siamese neural network bromley1993siamese ; burges .
Figure 4: Schematic picture of the model output combined with that of the next time step to form a loss consistent with a siamese neural network where the inputs of the two time steps are processed with the same model which is of the form of (10a). We considered the case of 3 input vectors $\mathbf{v}^{(i)}_{t}$ of a single run $n$ whose label we omitted to avoid clutter. For the picture we took $n_{\text{hidden}}=3$ whereas in reality the $n_{\text{hidden}}$ is as in (13).
After experimenting with (10), it is found that the trained function $h(\mathbf{V}_{t})$, unlike the real $H$-functional, displays strongly oscillatory behavior upon increasing the time step. To mitigate this, we note that (6a) is an inequality which only requires that the $H$-functional does not increase as a function of time, not that it decreases. Therefore, we consider a generalization of the loss function of (10b) which is given by
$\begin{split}&\mathcal{L}^{\alpha}\left(\mathbf{V}^{[n]}_{t},\mathbf{V}^{[n]}_{t+1}\right)\\ =&\text{leaky-ReLu}\left(h\left(\mathbf{V}^{[n]}_{t}\right)-h\left(\mathbf{V}^{[n]}_{t+1}\right),\alpha\right)\\ &+\lambda\,\!\left(h\left(\mathbf{V}^{[n]}_{t+1}\right)-h\left(\mathbf{V}^{[n]}_{t}\right)\right)^{2}~~.\end{split}$ (11)
It follows that when $0\leq\alpha<1$, (11) makes sure that the function $h(\mathbf{V}_{t})$ is punished more for incorrectly outputting a lower value at a later time step than it is rewarded for correctly outputting a higher value at later times. This choice is not only consistent with (6a), which only imposes the former on the $H$-functional, but also turns out to reduce oscillation of $h(\mathbf{V}_{t+1})$. We optimize for a total number of epochs $N_{e}=120$ with the loss of (11) averaged over runs $n$ and time steps $t$ using the Adam optimizer kingma2017adammethodstochasticoptimization , choosing learning rate $4\cdot 10^{-4}$ and $\lambda=0.2$. All the hidden layers of the model have hidden dimension $n_{\text{hidden}}$ and ReLu-activations (see Fig. 4 and Table 1). While training, we increase $\alpha$ as a function of epoch $e$ as
$\alpha(e)=1-\frac{1-\alpha_{\text{final}}}{N_{e}}\cdot e~~.$ (12)
We train three distinct models with hidden dimensions
$n_{\text{hidden}}\in\{16,32,64\}~~.$ (13)
Because of the choice of a gradual increase according to (12) with a high $\alpha_{\text{final}}$, one eventually reaches a point for all of the models where $\mathcal{L}_{t}^{\alpha(e)}>0$, which indicates that the trivial solution where $h(\mathbf{V}_{t+1})-h(\mathbf{V}_{t})=0$ for all $t$ is no longer suboptimal. This means the model can no longer improve and the model output difference between time steps starts to gradually move to zero. We call the epoch for which this first happens the critical epoch $e_{\text{c}}$. Table (2) gives the average loss for the different models at the final epoch together with $e_{\text{c}}$.
Layer Hyperparameter Activation
Input Input units: 2
Dense Output units: $n_{\text{hidden}}$ ReLU
Dense Output units: $n_{\text{hidden}}$ ReLU
Sum pooling
Dense Output units: $n_{\text{hidden}}$ ReLU
Dense Output units: 1 Linear
Table 1: Model architecture based on the Deepsets zaheer2018deepsets architecture described by (10a).
$n_{\text{hidden}}$ $\langle\mathcal{L}^{\alpha_{\text{final}}}(N_{e})\rangle$ $\langle\mathcal{L}^{(0)}(N_{e})\rangle$ $e_{\text{c}}$
16 0.00000950 -0.00001418 95
32 0.00017283 -0.00111001 105
64 0.00003145 -0.00038164 111
Table 2: Average loss of (11) at final epoch for the different models and the epoch $e_{\text{c}}$ for which $\langle\mathcal{L}^{\alpha(e)}\rangle$ starts to flip sign. We also provided the final loss with $\alpha=1$ of (10b) as this loss puts the different models on equal footing.
V Comparison with $H$-functional
Having fully specified the learning process of the model $h(\mathbf{V}_{t})$, let us consider its relation to the $H$-functional provided in (5). The model $h(\mathbf{V}_{t})$ can only possibly resemble the $H$-functional up to an affine transformation kuroyanagi2025deeplearningthermodynamiclaws ; rahaman2019learningarrowtime . This is because (11) is invariant under shifts constant in time. Furthermore, we have for $c\in\mathbb{R}$
$\{h(\mathbf{V}_{t}),\lambda\}\rightarrow\{ch(\mathbf{V}_{t}),\lambda/c\}\Rightarrow\mathcal{L}^{\alpha}_{t}\rightarrow c\mathcal{L}^{\alpha}_{t}~~,$ (14)
which means we can multiply the function $h(\mathbf{V}_{t})$ with a constant factor and it will only affect the loss with a total factor provided we rescale the physically meaningless $\lambda$. Therefore, in order to compare the model to the real $H$-functional we transform the function $h(\mathbf{V}_{t})$ as
$\tilde{h}^{[n]}\left(\mathbf{V}^{[n]}_{t}\right)=a^{[n]}h\left(\mathbf{V}^{[n]}_{t}\right)+b^{[n]}~~.$ (15)
The coefficients $a^{[n]}$ and $b^{[n]}$ are fitted to the $H$-functional with the least squares algorithm for each separate run $n$ and for all time steps $t$, i.e. they are computed as
$\operatorname*{argmin}_{a^{[n]},b^{[n]}}=\sum_{t}\left(\tilde{h}^{[n]}\left(\mathbf{V}^{[n]}_{t}\right)-H_{t}^{[n]}\right)^{2}~~.$ (16)
In Fig. 5, the result is shown for the different models and compared to the $H$-functional. It is clear that for $n_{\text{hidden}}=32$ the model performs better than for $n_{\text{hidden}}=16$, whereas there is no big difference between $n_{\text{hidden}}=64$ ad $n_{\text{hidden}}=32$. Although the affine parameters are fitted for each run separately, we would ideally have a model that can provide a value for the $H$-functional that is consistent throughout all runs. To find out whether the model is consistent throughout runs, we compute the average and the variance of the affine parameters in Table 3. We see that going from $n_{\text{hidden}}=16$ to $n_{\text{hidden}}=32$ lowers the error, which indicates a better learning process, whereas $n_{\text{hidden}}=64$ slightly increases it, although the coefficient of variation of the affine parameters is lowered significantly, suggesting that it is a more coherent model. Note that $\langle a\rangle$ is always negative as the model learns a function that increases in time whereas the $H$-functional is defined to be decreasing in time as follows from (6a).
$n_{\text{hidden}}$ MSE $\langle a\rangle$ $\langle b\rangle$ $\mathrm{CV}(a)$ $\mathrm{CV}(b)$
16 0.0429 -448.216 -9452.722 0.0430 0.0433
32 0.0155 -4.622 1104.860 0.0487 0.0462
64 0.0171 -14.885 -380.249 0.0215 0.0221
Table 3: Mean affine parameters $\langle a\rangle$, $\langle b\rangle$; coefficient of variation $\mathrm{CV}(x)=\sqrt{\mathrm{Var}(x)}/|\langle x\rangle|$; and mean-squared-error loss of the affine fits for the different models.
Figure 5: Picture of $H$-functional given by (5) vs affinely fitted model output. When computing the $H$-functional, we used the same bin size that was also used in Fig. 3 to perform a Riemann sum.
VI Discussion
In this work we used a siamese neural network to learn the $H$-functional of $H$-theorem for sequential data of a numerical simulation of a thousand hard disks colliding with each other after being initialized in a low-entropy state. The loss function that was formulated in order to learn $H$-theorem was based on the requirement that the learned function should attribute higher values to input data corresponding to later points in time. This requirement forces the model to identify the arrow of time which according to $H$-theorem was predicted to exist for this simple but also infinitely chaotic model of hard disks. The resulting model only resembles the extracted $H$-functional after an appropriately fitted affine transformation. Two things that were found to be important for making the model resemble the $H$-functional are
  1. 1.
    The DeepSets architecture zaheer2018deepsets , which implements the permutation invariance of the particle labels, thereby drastically simplifying the model compared to a model that treats every particle separately.
  2. 2.
    The sign-dependent tilt of the loss which makes it so that at late epochs, the model is penalized more strongly for outputting a decrease in time than it is rewarded for outputting an increase, thereby generating a model that is less oscillatory at late times and steeper at short times.
In future work, it would be interesting to explore more complicated types of collisional interactions, such as for example polyatomic gases Cercignani1981 or for active gas of two species of hard spheres that collide reactively and can resemble the phenomenon of bird flocking rubenliermicroscopic .
VII Methods
The training of the PyTorch model was performed on the Snellius supercomputer at SURF, the Dutch national supercomputing facility. The source code developed for this study is openly available at https://github.com/rubenlier/Machine-Learning-H-theorem.
VIII Acknowledgements
We thank Roy Stegeman and Karthik Viswanathan for useful discussions.